# Co-NAML-LSTUR: A Combined Model with Attentive Multi-View Learning and Long- and Short-term User Representations for News Recommendation

### v1

@@ Abstract
News recommendation systems play a vital role in mitigating information overload by delivering personalized news content. A central challenge is to effectively model both multi-view news representations and the dynamic nature of user interests, which often span both short- and long-term preferences. Existing methods typically rely on single-view features of news articles (e.g., titles or categories) or fail to comprehensively capture user preferences across time scales. In this work, we propose Co-NAML-LSTUR, a hybrid news recommendation framework that integrates NAML for attentive multi-view news modeling and LSTUR for capturing both long- and short-term user representations. Our model also incorporates BERT-based word embeddings to enhance semantic feature extraction. We evaluate Co-NAML-LSTUR on two widely used benchmarks, MIND-small and MIND-large. Experimental results show that Co-NAML-LSTUR achieves substantial improvements over most state-of-the-art baselines on MIND-small and MIND-large, respectively. These results demonstrate the effectiveness of combining multi-view news representations with dual-scale user modeling. The implementation of our model is publicly available at https://github.com/MinhNguyenDS/Co-NAML-LSTUR.
@@ I. Introduction
Motivated by these limitations, we propose Co-NAML-LSTUR (a Combined model with (Neural) Attentive Multi-view Learning and Long- and Short-term User Representations), a novel framework that combines the strengths of two state-of-the-art models, NAML and LSTUR. Co-NAML-LSTUR is designed to simultaneously learn multi-view news representations (titles, categories, abstracts) as in NAML, while also modeling both short-term behaviors (recently read articles) and long-term user preferences (historical clicks over time) as in LSTUR. Furthermore, inspired by DKN, we introduce a neural network-based scoring function for candidate news matching. To improve semantic understanding of news content, we also leverage BERT-base word embeddings to enhance the quality of text representations.
The main contributions of this work can be summarized as follows:
  • We propose Co-NAML-LSTUR, a hybrid model that integrates attentive multi-view news representation and long- and short-term user modeling for personalized news recommendation.
  • We design and evaluate several model variants to study the effectiveness of different components within the Co-NAML-LSTUR framework.
  • We conduct extensive experiments on the MIND dataset (including MIND-small and MIND-large) and demonstrate that our model achieves competitive or superior performance compared to strong baselines.
@@ 2.3 Neural News Recommendation
Traditional ID-based methods often struggle with the cold-start problem in news recommendation due to the rapid update and short lifespan of news articles (Wu et al. 2020). To overcome these issues, neural network-based methods have been proposed to automatically learn representations of news and users (Okura et al. 2017). Sequential recommendation approaches have also gained attention, aiming to model user behaviors with deep learning-based news recommendation methods, such as RNNs, CNNs, or self-attention networks (Wu et al. 2019b; An et al. 2019c; Khattar et al. 2018; Kumar et al. 2017). For instance, (Okura et al. 2017) proposed to learn news embeddings from article bodies using denoising autoencoders and capture user interests using GRU networks. (Wang et al. 2018) introduced a knowledge-aware CNN that incorporates information from knowledge graphs to encode news titles.
@@
III. MINDtiny
3.1 Overview of the MIND Dataset
Statistic MIND-small MIND-large
#News 65,238 161,013
#Categories 18 20
#Impressions 230,117 15,777,377
#Clicks 347,727 24,155,470
3.2 Motivation and Construction of MINDtiny
While the full-scale MIND dataset is highly comprehensive, its size poses significant challenges for training on resource-constrained devices. To overcome this limitation while retaining the essential properties of large-scale user behavior and linguistic diversity, we introduce MINDtiny—a compact yet representative dataset tailored for efficient model training and experimentation. The construction of MINDtiny is guided by two key objectives: (1) to preserve temporal behavioral patterns of users, and (2) to retain representative and frequently interacted news content. To achieve this, we apply a selective sampling strategy from MIND-large based on the following criteria:
By combining these filters, MINDtiny preserves the essential behavioral and semantic characteristics of MIND-large while significantly reducing the overall data size. In our experiments, MINDtiny is used in conjunction with MIND-small to form a compact yet informative training set for our proposed model. This combined dataset enables efficient training in low-resource settings while still capturing the diversity and complexity necessary for modeling realistic user preferences and news content. This strategy allows us to balance computational efficiency with the retention of critical information patterns from the original large-scale MIND dataset, making it well-suited for both model development and robust evaluation.
3.3 Dataset Statistics and Analysis
Detailed statistics of MINDtiny are presented in Table 2. The dataset comprises 4,844 users and 19,557 unique news articles. It includes 16 top-level categories and 23 subcategories. On average, the abstract length is approximately 30 words, while titles are concise, with an average of 10.88 words.
Statistic Value
#News 19,557
#Users 4,844
#Categories 16
#Subcategories 23
Avg. Title Length 10.88 words
Avg. Abstract Length 30.02 words
Table 2: Statistics of the MINDtiny dataset.
Figure 3: Distributional statistics of categories, title lengths, and abstract lengths in MINDtiny.
IV. Our Method
@@ 4.1 News Encoder
Category Encoder. Many news platforms tag articles with hierarchical topic labels (e.g., categories and subcategories), which provide high-level semantic cues about the content. To incorporate this metadata, we encode both category ID $v_{c}$ and subcategory ID $v_{sc}$ via dense embeddings, initialized using DistilBERT and further transformed through nonlinear projections:
@@ 4.2 User Encoder
where $H(\cdot)$ is a feedforward neural network (DNN), and [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes vector concatenation. The final attention-based user embedding $u_{att}$ is computed as the weighted sum of historical embeddings:
@@ 4.4 Model Training
Following prior work (Wu et al. 2019a), we adopt the Noise Contrastive Estimation (NCE) loss for model training. For each news article that has been clicked by a user (treated as a positive sample, ${{\widehat{y}}_{i}^{+}}$), we randomly sample $K$ news articles that were not clicked by the same user (treated as negative samples, $[{\widehat{y}}_{1}^{-},{\widehat{y}}_{2}^{-},\dots,{\widehat{y}}_{K}^{-}]$). We then predict the click probabilities for the entire set $S$ consisting of both positive and negative samples.
@@ 5.1 Experiment Setup
We conduct our experiments using the MIND dataset (Wu et al. 2020), a real-world news recommendation dataset collected from user behavior logs on Microsoft News. It is available in two versions: MIND-large, which contains over 15 million impression logs generated by 1 million users, and MIND-small, a subset randomly sampled from 50,000 users, as detailed in Table 1. During training, we combine MIND-small and MIND-tiny to form the training set. Evaluations are conducted separately on MIND-small and MIND-large to assess generalizability across different data scales. See Section III.MINDtiny for a summary of dataset statistics.
@@ 5.3 Main Results
On the MIND-small benchmark, Co-NAML-LSTUR achieves an AUC of 0.6571 and an MRR of 0.3119, surpassing several strong baselines such as NRMS (+4.2% AUC), LSTUR (+2.0% MRR), and NAML (+4.0% nDCG@10), highlighting its effective modeling of both semantic and temporal user signals. While MINER yields the best overall performance, Co-NAML-LSTUR attains the second-highest scores across all metrics, despite using a relatively lightweight architecture (46.4M parameters compared to LSTUR’s 71.6M).
@@
Comparison with state-of-the-art models.
Although MINER reports the highest performance on both benchmarks, it requires modeling multiple user interest vectors and computing adaptive selection weights, which increases inference complexity. In contrast, Co-NAML-LSTUR achieves a competitive trade-off between accuracy and computational cost, making it more scalable in real-world deployments.
@@ 5.5 Effectiveness of Click Prediction
Table 5: Effect of click prediction module design on recommendation performance. The neural predictor consistently outperforms the dot-product variant across all evaluation metrics.
@@ VI. Conclusion
In this paper, we propose Co-NAML-LSTUR, a unified and modular neural framework for personalized news recommendation. Our approach addresses key challenges in modeling both user interest diversity and the semantic richness of news content. To represent news articles, we design a hybrid encoder that integrates category-aware and content-sensitive features through attentive multi-view learning. On the user side, we fuse long- and short-term user preferences by leveraging the strengths of both static profile embeddings and sequential browsing behavior modeling, allowing the model to adapt to dynamic reading patterns. A central component of our framework is the incorporation of a click prediction module, which can flexibly switch between dot-product and neural interaction functions. Our experiments demonstrate that the neural variant significantly enhances recommendation accuracy, albeit with increased computational cost. We conduct extensive evaluations on the MIND dataset, showing that Co-NAML-LSTUR outperforms several state-of-the-art baselines across multiple ranking metrics. Additionally, our case study and visual analysis reveal that the model is capable of capturing multi-faceted user interests and generating relevant, diverse recommendations. The modularity of Co-NAML-LSTUR makes it a promising foundation for future work. It can be easily extended with external knowledge, multi-modal content, or user demographic features to further boost performance and interpretability.
Acknowledgements
We sincerely thank Microsoft News for granting access to the MIND dataset. We also gratefully acknowledge the open-source community, particularly the maintainers of the https://github.com/yusanshi/news-recommendation, whose implementations of several baseline models significantly contributed to the development and evaluation of our framework. Including NRMS (Neural News Recommendation with Multi-Head Self-Attention), NAML (Neural News Recommendation with Attentive Multi-View Learning), LSTUR (Neural News Recommendation with Long- and Short-Term User Representations), DKN (Deep Knowledge-Aware Network), Hi-Fi Ark (High-Fidelity User Representation), and TANR (Topic-Aware Neural News Recommendation). Their efforts have greatly facilitated reproducible and extensible research in the news recommendation domain.

### v2

@@ Abstract
News recommendation systems play a critical role in alleviating information overload by delivering personalized content. A key challenge lies in jointly modeling multi-view representations of news articles and capturing the dynamic, dual-scale nature of user interests-encompassing both short- and long-term preferences. Prior methods often rely on single-view features or insufficiently model user behavior across time. In this work, we introduce Co-NAML-LSTUR, a hybrid news recommendation framework that integrates NAML for attentive multi-view news encoding and LSTUR for hierarchical user modeling, designed for training on limited data resources. Our approach leverages BERT-based embeddings to enhance semantic representation. We evaluate Co-NAML-LSTUR on two widely used benchmarks, MIND-small and MIND-large. Results show that our model significantly outperforms strong baselines, achieving improvements over NRMS by 1.55% in AUC and 1.15% in MRR, and over NAML by 2.45% in AUC and 1.71% in MRR. These findings highlight the effectiveness of our efficiency-focused hybrid model, which combines multi-view news modeling with dual-scale user representations for practical, resource-limited resources rather than a claim to absolute state-of-the-art (SOTA). Code is publicly available. The implementation of our model is publicly available at https://github.com/MinhNguyenDS/Co-NAML-LSTUR.
@@ I. Introduction
Motivated by limitations, we propose Co-NAML-LSTUR (a Combined model with (Neural) Attentive Multi-view Learning and Long- and Short-term User Representations), a novel framework that combines the strengths of two state-of-the-art models. Co-NAML-LSTUR is designed to simultaneously learn multi-view news representations (titles, categories, abstracts), while also modeling both short-term behaviors (recently read articles) and long-term user preferences (historical clicks over time). Furthermore, we use a neural network-based scoring function for candidate news matching and leverage BERT-base word embeddings to enhance the quality of text representations. The main contributions can be summarized as follows:
  • We propose Co-NAML-LSTUR, a hybrid model that integrates attentive multi-view news representation with long- and short-term user modeling for personalized news recommendation, designed for training with fewer resources compared to state-of-the-art methods.
  • We introduce MIND-tiny, a compact yet subset of MIND dataset designed to enable efficient model training on resource-constrained devices.
  • Experimental results on the MIND show that Co-NAML-LSTUR achieves competitive performance compared to strong baselines, confirming the effectiveness of our dual-view encoder and attention-based user modeling.
@@ 2.3 Neural News Recommendation
Traditional ID-based methods often struggle with the cold-start problem in news recommendation due to the rapid update and short lifespan of news articles (Wu et al. 2020). To overcome these issues, neural network-based methods have been proposed to automatically learn representations of news and users (Okura et al. 2017). Sequential recommendation approaches have also gained attention, aiming to model user behaviors with deep learning-based news recommendation methods, such as RNNs, CNNs, or self-attention networks (Wu et al. 2019b; An et al. 2019c; Nguyen et al. 2024; Khattar et al. 2018; Kumar et al. 2017). For instance, (Okura et al. 2017) proposed to learn news embeddings from article bodies using denoising autoencoders and capture user interests using GRU networks. (Wang et al. 2018) introduced a knowledge-aware CNN that incorporates information from knowledge graphs to encode news titles.
@@ 3.1 News Encoder
Category Encoder. Many news platforms tag articles with hierarchical topic labels (e.g., categories and subcategories), which provide high-level semantic cues about the content. To incorporate this metadata, we encode both category ID $v_{c}$ and subcategory ID $v_{sc}$ via dense embeddings, initialized using DistilBERT and nonlinear projections:
@@ 3.2 User Encoder
where $H(\cdot)$ is a feedforward neural network (DNN), and $\cdot$$\cdot$] denotes vector concatenation. The final attention-based user embedding $u_{att}$ is computed as the weighted sum of historical embeddings:
@@ 3.4 Model Training
Following prior work (Wu et al. 2019a), we adopt the Noise Contrastive Estimation (NCE) loss for model training. For each news article that has been clicked by a user (treated as a positive sample, ${\widehat{y}}_{i}^{+}$}), we randomly sample $K$ news articles that were not clicked by the same user (treated as negative samples, ${\widehat{y}}_{1}^{-},{\widehat{y}}_{2}^{-},\dots,{\widehat{y}}_{K}^{-}]$). We then predict the click probabilities for the entire set $S$ consisting of both positive and negative samples.
@@
IV. MIND-tiny
4.1 Overview of the MIND Dataset
Statistic MIND-small MIND-large
#News 65,238 161,013
#Categories 18 20
#Impressions 230,117 15,777,377
#Clicks 347,727 24,155,470
4.2 Motivation and Construction of MIND-tiny
While the full-scale MIND dataset is highly comprehensive, its size poses significant challenges for training on resource-constrained devices. To overcome this limitation while retaining the essential properties of large-scale user behavior and linguistic diversity, we introduce MIND-tiny—a compact yet representative dataset tailored for efficient model training and experimentation. The construction of MIND-tiny is guided by two key objectives: (1) to preserve temporal behavioral patterns of users, and (2) to retain representative and frequently interacted news content. To achieve this, we apply a selective sampling strategy from MIND-large based on the following criteria:
By combining these filters, MIND-tiny preserves the essential behavioral and semantic characteristics of MIND-large while significantly reducing the overall data size. This strategy allows us to balance computational efficiency with the retention of critical information patterns from the original large-scale MIND dataset, making it well-suited for training model. [^3] Filter high-activity users may underrepresent casual readers, potentially biasing results and limiting generalizability; future work should balance efficiency with broader audience diversity.
4.3 Dataset Statistics and Analysis
Detailed statistics of MIND-tiny are presented in Table 2. The dataset comprises 4,844 users and 19,557 unique news articles. It includes 16 top-level categories and 23 subcategories. On average, the abstract length is approximately 30 words, while titles are concise, with an average of 10.88 words.
Statistic Value
#News 19,557
#Users 4,844
#Categories 16
#Subcategories 23
Avg. Title Length 10.88 words
Avg. Abstract Length 30.02 words
Table 2: Statistics of the MIND-tiny dataset.
Figure 4: Distributional statistics of categories, title lengths, and abstract lengths in MIND-tiny.
@@ 5.1 Experiment Setup
We conduct our experiments using the MIND dataset (Wu et al. 2020), a real-world news recommendation dataset collected from user behavior logs on Microsoft News. It is available in two versions: MIND-large, which contains over 15 million impression logs generated by 1 million users, and MIND-small, a subset randomly sampled from 50,000 users, as detailed in Table 1. During training, we combine MIND-small and MIND-tiny to form the training set. Evaluations are conducted separately on MIND-small and MIND-large to assess generalizability across different data scales. See Section III.MIND-tiny for a summary of dataset statistics.
@@ 5.3 Main Results
On the MIND-small dataset, Co-NAML-LSTUR achieves an AUC of 0.6571 and an MRR of 0.3119, surpassing several strong baselines such as NRMS (+4.2% AUC), LSTUR (+2.0% MRR), and NAML (+4.0% nDCG@10), highlighting its effective modeling of both semantic and temporal user signals. While MINER yields the best overall performance, Co-NAML-LSTUR attains the second-highest scores across all metrics.
@@
Comparison with state-of-the-art models. Although MINER reports the highest performance on both benchmarks, it requires modeling multiple user interest vectors and computing adaptive selection weights, which leads to greater processing overhead. In contrast, Co-NAML-LSTUR achieves a competitive balance between accuracy and processing efficiency; our model—although trained on fewer resources by concatenating data from MIND-small and MIND-tiny—still delivers performance comparable to MINER, which is trained on the much larger MIND-large dataset. Overall, these results validate the core hypothesis of this work: that incorporating multi-view news semantics with an attentive long- and short-term user modeling pipeline yields meaningful performance benefits.
@@ 5.5 Effectiveness of Click Prediction
Table 5: Effect of click prediction module design on recommendation performance. The neural predictor consistently outperforms the dot-product across all evaluation metrics.
@@ VI. Conclusion
In this paper, we propose Co-NAML-LSTUR, a unified and modular neural framework for personalized news recommendation. Our approach addresses key challenges in modeling both user interest diversity and the semantic richness of news content. To represent news articles, we design a hybrid encoder that integrates category-aware and content-sensitive features through attentive multi-view learning. On the user side, we fuse long- and short-term user preferences by leveraging the strengths of both static profile embeddings and sequential browsing behavior modeling, allowing the model to adapt to dynamic reading patterns. A central component of our framework is the incorporation of a click prediction module, which can flexibly switch between dot-product and neural interaction functions. Our experiments demonstrate that the neural variant significantly enhances recommendation accuracy, albeit with increased computational cost. We conduct extensive evaluations on the MIND dataset, showing that Co-NAML-LSTUR outperforms several state-of-the-art baselines across multiple ranking metrics. Additionally, our case study and visual analysis reveal that the model is capable of capturing multi-faceted user interests and generating relevant, diverse recommendations. Our method makes it a promising foundation for future work. It can be extended with external knowledge, multi-modal content, or user demographic features to further boost performance and interpretability (such as analyze attention weight distributions, feature importance, and user group variations, provides quantitative interpretability or user-behavior analysis).
Authorship contribution statement.
Minh Hoang Nguyen: Methodology, Conceptualization, Writing – structuring original draft, Writing—reviewing. Thuat Thien Nguyen: Preprocessing data, Building dataset, Conceptualization, Writing – structuring original draft. Minh Nhat Ta: Preprocessing data, Writing – structuring original draft. Tung Le: Supervision, Supporting, Reviewing and Editing. Huy Tien Nguyen: Supervision, Supporting, Conceptualization, Project Administration.
Acknowledgements.
This research is partially funded by the Vingroup Innovation Foundation (VINIF) under the grant number VINIF.2021.JM01.N2
Disclosure of Interests.
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.